give an example of a non–trivial homomorphism or explain why none exists

φ:S4 → S3

1
Expert's answer
2021-05-07T09:28:54-0400

An example could by a homomorphism that is composed by two homomorphisms:

  1. There is a homomorphism "signature" "\text{sgn}: S_n \to \mathbb{Z}/2\mathbb{Z}" that associates to a permutation its signature.
  2. There is an injection "\iota: \mathbb{Z}/2\mathbb{Z} \to S_m" for any "m\geq 2" that associates "1" to a permutation "\begin{pmatrix} 1 & 2 \end{pmatrix}".

Composition of these two homomorphisms works, in particular, for "n=4, m=3" which gives us a homomorphism "\phi: S_4 \to S_3" with "\phi(even)=id; \phi(odd)=\begin{pmatrix} 1 & 2 \end{pmatrix}".


There exist another non-trivial homomorphism, that is more interesting because it is surjective. We can prove that "K=\{id, (12)(34), (13)(24), (14)(23) \}" is a normal subgroup of "S_4" and the quotient "S_4/K \simeq S_3", so the composition of these two surjective homomorphisms is a surjective homomorphism.